OpenStudy (anonymous):

solve the system by either elimination or substitution 3x-5y=21 2x+y=1 A.2,-3 B.2,3 C.-2,-1 D.-2,1 My guess is c

OpenStudy (freckles):

you could multiply the second equation by 5 and then add the equations together afterwards thus eliminating the y

OpenStudy (freckles):

what would the second equation look like after multiplying it by 5?

OpenStudy (freckles):

\[2x+y=1 \\ 5(2x+y)=5(1) \\ \text{ use distributive property on the left } \\ 5(2x+y)=5\]

OpenStudy (freckles):

remember that is a(b+c)=ab+ac this is distributive property

OpenStudy (anonymous):

i got 10x+5y=5

OpenStudy (freckles):

 ok add the equations now